Determine whether the distribution represents a probability distribution. Explain your answer.

X

1

3

5

7

P(X)

0.35

0.25

0.22

0.12


1
Expert's answer
2022-03-10T18:37:38-0500

For this distribution to be a probability distribution it is required that the sum of probabilities must be 1

"P(X=1)+P(X=3)+P(X=5)+P(X=9)=0.35+0.25+0.22+0.12=0.94\not=1"

So, it is not a probability distribution